# references.bib

### v1

@@ 7.3 Applying Our Recommendations to Apple Intelligence
In Case I, no user data is used for training [appledeviceandservermodels], which is compliant with Recommendation 1. Further, all processing of queries—including, importantly, queries involving E2EE content—occurs locally on device [applepressrelease]. Therefore, Case I is also fully compliant with E2EE security according to Recommendation 2, and features relying on Case I may be publicized as E2EE without qualifications according to Recommendation 3. Notably, Apple Intelligence prioritizes Case I whenever possible, as specified in Recommendation 2(a). Lastly, at the time of writing, Apple Intelligence is off by default, and has to be manually activated by users in their device settings [arstechnica-apple-intelligence-off-default], which is in line with Recommendation 4.
@@
Just like Case I, no user data is used to train Apple’s PCC models [appledeviceandservermodels], which is compliant with Recommendation 1. However, notwithstanding the security-conscious design choices within Apple’s PCC, processing inference queries inside trusted hardware provides a different kind of security from E2EE security, and thus is not compatible with E2EE under Recommendation 2. Therefore, in line with Recommendation 3, Apple should not advertise Case II as E2EE without qualifications; rather, Apple should clearly articulate the limitations introduced by the TEE processing, and explicitly state the contexts in which E2EE applications (both native and third-party) no longer meet E2EE guarantees. Such notices should be prominently displayed. Importantly, even users who have Apple Intelligence off should be aware of these limitations, due to the systemic E2EE properties mentioned in Section 2.1.2. Finally, Case II appears to adhere to Recommendation 4, as Apple Intelligence as a whole is off by default and, to our limited knowledge, seems to have reasonable opt-in consent mechanisms.
@@
On Recommendation 1, if users do not sign in to their ChatGPT accounts, then OpenAI does not use requests to train its models [applesupport-chatgptext]—thus, Case III processing is compliant with Recommendation 1 for users who are not logged in. However, if a user is logged into their OpenAI account and has not configured their settings otherwise [^58] ChatGPT accounts have model training on by default [openai-datacontrols]., queries outsourced to ChatGPT may be used for training OpenAI models [openaisupport-appleintelligence-datahandling], directly opposing Recommendation 1. Thus, overall, Case III does not uphold Recommendation 1.
On Recommendation 2, if Siri is integrated within an application that is E2EE, then processing queries with ChatGPT via Siri would be in violation of Recommendation 2(b)(i), as E2EE content would be used by a third party. This is also true if ChatGPT can respond to requests on third-party E2EE or ADP-E2EE applications. Furthermore, if a user has logged into their OpenAI account and has not configured their settings otherwise, ChatGPT can log requests and use query data to train its models. Using ChatGPT to process queries on E2EE content would therefore be in violation of Recommendation 2(b)(ii), since the user’s E2EE content would not solely be used for the purpose of fulfilling that user’s requests.
Hence, according to Recommendation 3, Apple should not advertise Case III as E2EE without qualifications; it should clearly disclose the contexts in which E2EE data could be processed by third parties and how those third parties could use it.
Finally, since the ChatGPT extension in Siri is off by default and must be explicitly enabled in settings [applesupport-chatgptext], Case III abides by Recommendation 4. Furthermore, it seems users are always made aware of when ChatGPT is being used to respond to a query [applesupport-siri]. Even if a user chooses not to be asked for permission before each query is sent to OpenAI, a small banner remains at the bottom of the response stating that it has been generated via ChatGPT. An additional disclosure that content sent to ChatGPT is no longer E2EE could further clarify security guarantees to users.
@@ 8 Discussion
Law enforcement agencies and the global intelligence community have sought over decades to both limit the use of strong encryption and exploit the naturally occurring weaknesses of encrypted systems [bellovin2013wiretapping]. [^59] Very recent developments suggest an interesting possible shift, though it is too soon to be clear [CISA-FBI-2024, knodel-tpp24]. See also [baker19]. These efforts are in response to the technical reality that end-to-end encryption provides confidentiality to anyone who uses it. The so-called “Crypto Wars” can be described as the confrontation across regulatory, corporate, and research settings between state power and encryption proponents to determine whether there exists a technical compromise that would facilitate “lawful access” to confidential data [NAP25010].
The Crypto Wars have evolved as encryption has evolved and technical proposals have evolved from weak cryptographic standards to various forms of key escrow [abelson2015doormats]. The latest battles in the protracted war between lawful access and secure encryption are characterized by proposed technology solutions framed as sophisticated alternatives to backdoors. Rather than mandating the implementation of weak encryption protocols that allow a third-party to exfiltrate data, design features like client-side scanning would preserve strong encryption algorithms while accessing E2EE content in other ways, namely on the device (the “client”) and in the cloud using cryptographic computational techniques not requiring decryption.
What do backdoors, client-side scanning and homomorphic encryption have to do with AI assistants? Like data-driven algorithms, lawful access is incompatible with the privacy and confidentiality of end-to-end encryption, especially when users are uninformed of the risks. Furthermore, they are extensions or features of private messaging applications, not meant to enhance privacy and confidentiality; rather, they are designed to undermine it by accessing data that would not otherwise be accessible. A recurring concern in the Crypto Wars has been the risk of large-scale, hard-to-detect abuse and compromise: Even if we all agreed that AI assistants are beneficial, if E2EE data can be accessed by AI assistants for good, that method of access could also be used to track and target users, undermining the core of the security guarantee that E2EE provides.
@@ 8.2 Business Incentives to Integrate AI (and Fundamental Rights)
Companies with data-driven business models may see messaging data as a valuable resource for refining AI algorithms. These companies might be more inclined to collect and analyze user interactions to enhance personalization, improve accuracy, or monetize insights, than they are offer E2EE, which places strong limits on what the platform can do. The temptation to use this data for training AI models will likely grow as AI assistants become increasingly embedded in everyday devices. Adding to these incentives in the private messaging context is the fact that messaging data is a potentially valuable source of untapped human-authored data in an environment where concerns are arising about the scarcity of such data.
E2EE has become established as a key feature of private messaging and a key part of major messaging providers’ offerings and marketing. Interestingly, in this case, the business incentives to protect proprietary and user data appear aligned with human rights protections. However, within the last couple of years, we have seen a shift towards data-heavy AI features, following the public release of ChatGPT by OpenAI. In what appears to be the new business model for Apple—which for years built its brand around privacy—and its competitors, services are seeking out and leveraging large sources of user data. The speed with which companies are turning to AI tools, including incorporating them into E2EE environments, is happening at a pace that does not permit thorough community deliberation over associated risks and externalities; and the trend appears to be corporate-driven even in the absence of clear user demand, and in the face of user concerns [cicek2024adverse, kikuchi2023growing]. Yet informed debate and discussion to consider the compatibility of privacy tools like end-to-end encryption and AI, data-driven features requires human rights and legal advocates, as well as technologists, to carefully consider impacts and harms in a more deliberate process.
@@ 8.3 But Can’t I Just Paste That Into ChatGPT?
In a nutshell, though, the two key differences are in the service provider’s promise about an E2EE application—that it does not access or process E2EE content, that it designs its systems to uphold E2EE confidentiality—and in the systemic nature of the breach of confidentiality. If a user pastes their E2EE messages into an AI assistant, or indeed if they post them publicly online, that is a intentional act undertaken by an individual that may be a breach of social norms (depending on the context) but neither breaches the E2EE service provider’s promise or creates a systemic breach of confidentiality by which similar processes can happen as a matter of course, as part of a regular app feature that is easily accessible and whose frequent use is encouraged by the very service provider who offers the E2EE.
@@ 9 Conclusion
AI models are being developed at a rapid pace, and being incorporated into ubiquitous applications, including E2EE systems. In this work, we performed a detailed examination of the technical and legal tensions between AI models and E2EE applications, identifying serious concerns.
Deployment of AI features in E2EE environments should be accompanied by informed discussions on potential risks and harms. Our work represents a first step towards this; we hope this conversation continues involving free speech, privacy, and human rights advocates, civil society representatives, and others outside academia, to develop frameworks and standards for the responsible deployment of AI and E2EE. It is critical for these debates to occur promptly, before unilateral industry rollout.
@@
MK had the initial idea and co-led the overall structuring of the paper and recommendations with SP. AF, BLH, and DY contributed equally to designing and writing technical background, implications, and recommendations. DF and JL contributed equally to designing and writing legal background, implications, and recommendations. SA led the Apple case study and provided research support throughout other aspects of the paper. SP was faculty lead and KC provided consultation.
@@ Acknowledgments
Appendix A E2EE Features Evaluation Framework
@@
Our framework distinguishes between three core axes that must be considered in the context of application features: (1) the technical implementation of the feature, and its confidentiality guarantees (or lack thereof); (2) the data inputs to the feature, i.e., what type of user data will interface with it; and (3) the level of user control over the feature. We discuss these components in more detail below.

### v2

@@ 7.3 Applying Our Recommendations to Apple Intelligence
In Case I, no user data is used for training [appledeviceandservermodels], which is compliant with Recommendation 1. Further, all processing of queries—including, importantly, queries involving E2EE content—occurs locally on device [applepressrelease]. Therefore, Case I is also fully compliant with E2EE security according to Recommendation 2, and features relying on Case I may be publicized as E2EE without qualifications according to Recommendation 3. Apple claims to prioritize Case I; “many of the models” run fully on-device and it is the “more complex requests” that require using Cases II or III [applepressrelease]—though it is not clear exactly what this means, the statement appears to align with Recommendation 2(a). Lastly, during its first release in iOS 18.1, Apple Intelligence was off by default, and had to be manually activated by users in their device settings [applereleasesupport], which is in line with Recommendation 4. As of iOS 18.3, Apple Intelligence is on by default and “integrated across features in your apps” [apple-intelligence-ios183]. This is in violation of our Recommendation 4.
@@
Just like Case I, no user data is used to train Apple’s PCC models [appledeviceandservermodels], which is compliant with Recommendation 1. However, notwithstanding the security-conscious design choices within Apple’s PCC, processing inference queries inside trusted hardware provides a different kind of security from E2EE security, and thus is not compatible with E2EE under Recommendation 2. Therefore, in line with Recommendation 3, Apple should not advertise features relying on Case II as E2EE without qualifications; rather, Apple should clearly articulate the limitations introduced by the TEE processing, and explicitly state the contexts in which E2EE applications (both native and third-party) no longer meet E2EE guarantees. Such notices should be prominently displayed. Importantly, even users who have Apple Intelligence off should be aware of these limitations, due to the systemic E2EE properties mentioned in Section 2.1.3. Finally, during its first release in iOS 18.1, Apple Intelligence was off by default, and had to be manually activated by users [applereleasesupport], which is in line with Recommendation 4. However, as of iOS 18.3, Case II no longer adheres to Recommendation 4, as Apple Intelligence is on by default [apple-intelligence-ios183].
@@
On Recommendation 1, if users do not sign in to their ChatGPT accounts, then OpenAI does not use user queries to train its models [applesupport-chatgptext]—thus, Case III processing is compliant with Recommendation 1 for users who are not logged in. However, if a user is logged into their OpenAI account and has not configured their settings otherwise [^72] ChatGPT accounts have model training on by default [openai-datacontrols]., queries outsourced to ChatGPT may be used for training OpenAI models [openaisupport-appleintelligence-datahandling], directly opposing Recommendation 1. Thus, overall, Case III does not uphold Recommendation 1.
On Recommendation 2, if Siri is integrated within an application that is E2EE, then processing queries with ChatGPT via Siri would be in violation of Recommendation 2(b)(i), as E2EE content would be used by a third party. This is also true if ChatGPT can respond to queries involving third-party E2EE or ADP-E2EE application content. Furthermore, if a user has logged into their OpenAI account and has not configured their settings otherwise, ChatGPT can log requests and use query data to train its models. Using ChatGPT to process queries on E2EE content would therefore be in violation of Recommendation 2(b)(ii), since the user’s E2EE content would not solely be used for the purpose of fulfilling that user’s requests.
Hence, according to Recommendation 3, Apple should not advertise features relying on Case III as E2EE without qualifications; it should clearly disclose the contexts in which E2EE data could be processed by third parties and how those third parties could use it.
Finally, since the ChatGPT extension in Siri is off by default and must be explicitly enabled in settings [applesupport-chatgptext], Case III abides by Recommendation 4. Furthermore, it seems users are always made aware of when ChatGPT is being used to respond to a query [applesupport-siri]. Even if a user chooses not to be asked for permission before each query is sent to OpenAI, a small banner remains at the bottom of the response stating that it has been generated via ChatGPT. [^73] For an example of what this looks like, see this example from the Apple Intelligence press release [applepressrelease]. An additional disclosure that content sent to ChatGPT is no longer E2EE could further clarify security guarantees to users.
@@ 8 Discussion
Law enforcement agencies and the global intelligence community have sought over decades to both limit the use of strong encryption and exploit the naturally occurring weaknesses of encrypted systems  [^74] Very recent developments suggest an interesting possible shift, though it is too soon to be clear [CISA-FBI-2024, knodel-tpp24]. See also [baker19]. [bellovin2013wiretapping]. These efforts are in response to the technical reality that end-to-end encryption provides confidentiality to anyone who uses it, including criminals. The so-called “Crypto Wars” can be described as the confrontation across regulatory, corporate, and research settings between state power and encryption proponents to determine whether there exists a technical compromise that would facilitate “lawful access” to confidential data [NAP25010]. The Crypto Wars have evolved over the years with encryption, and technical proposals have evolved from pushing for weak cryptographic standards to various forms of key escrow [abelson2015doormats]. The latest proposed technological solutions are framed as sophisticated alternatives to backdoors, which would preserve strong encryption algorithms while accessing E2EE content in other ways, namely on the device (the “client”) and in the cloud using cryptographic computational techniques not requiring decryption.
What do backdoors, client-side scanning and homomorphic encryption have to do with AI assistants? Like data-driven algorithms, lawful access is incompatible with the privacy and confidentiality of end-to-end encryption, especially when users are uninformed of the risks. Furthermore, both are extensions or features of private messaging applications, not meant to enhance privacy and confidentiality; rather, they are designed to undermine it by enabling access to data that would not otherwise be accessible. A recurring concern in the Crypto Wars has been the risk of large-scale, hard-to-detect abuse and compromise: Even if we all agreed that AI assistants are beneficial, if E2EE data can be accessed by AI assistants for good, that method of access could also be used to track and target users, undermining the core of the security guarantee that E2EE provides.
@@ 8.2 Business Incentives to Integrate AI (and Fundamental Rights)
Companies with data-driven business models may see messaging data as a valuable resource for refining AI algorithms, and the temptation to use this data will likely grow as AI assistants become increasingly embedded in everyday devices. Messaging data is a source of largely untapped human-authored text at a time when concerns are mounting about the scarcity of such data; providing E2EE hinders the platform’s ability to collect and analyze this data.
We are concerned that this new business incentive may eclipse the existing business incentives to protect user data, which are better aligned aligned with human rights protections. In what appears to be the new business model for Apple—which for years built its brand around privacy—and its competitors, services are seeking out and leveraging large sources of user data at a pace that does not permit thorough community deliberation over associated risks and externalities, even in the absence of clear user demand, and in the face of user concerns [cicek2024adverse, kikuchi2023growing].
@@ 8.3 But Can’t I Just Paste That Into ChatGPT?
In a nutshell, the two key differences are in the service provider’s promise about an E2EE application—that it does not access or process E2EE content, that it designs its systems to uphold E2EE confidentiality—and in the systemic nature of the breach of confidentiality. If a user pastes their E2EE messages into an AI assistant, or indeed if they post them publicly online, that is a intentional act undertaken by an individual that may be a breach of social norms (depending on the context) but neither breaches the E2EE service provider’s promise or creates a systemic breach of confidentiality by which similar processes can happen as a matter of course, as part of a regular app feature that is easily accessible and whose frequent use is encouraged by the very service provider who offers the E2EE.
8.4 Additional Security Concerns: Message Integrity
This work primarily focuses on the interplay between AI assistants and the confidentiality and systemic guarantees of E2EE. Yet, as discussed in Section 2.1.2, modern E2EE applications aim to provide additional security properties beyond these two. An important direction for future work is to perform an analysis of the (in)compatibility of AI assistants and these other guarantees. For any security property that an E2EE application claims to provide, it is critical that the implementation of AI assistants upholds this guarantee.
Integrity is another standard security guarantee of E2EE applications, which states that a message received by the recipient should be equal to the one sent by the sender. AI summary features could be in tension with integrity: incorrect summaries of messages could be considered a violation of this guarantee, as they display an inaccurate representation of the message to the recipient. While the original (unmodified) messages are still visible in the chat history, incorrect summaries may still cause unintended consequences akin to if the messages were modified in transit. Incorrect summaries may occur accidentally, as a result of imperfect AI features, or deliberately, as a result of targeted attacks. For example, consider a group chat with Alice, Bob, and Eve, where AI summaries are enabled. If Alice is idle for a period of time while Bob and Eve send messages, and if Eve can trick the AI feature into misrepresenting Bob’s messages in the summary, this could create confusion or conflict between Alice and Bob.
Whether and in what contexts incorrect AI summaries could or should be treated as a violation of integrity or not is an interesting question, outside the scope of this work. The analysis is more subtle than for confidentiality, since integrity violations can in principle be detected by users by inspecting the original messages in the full chat history—in contrast, once E2EE content has been disclosed to third parties, it is not possible to backtrack. Exploring this question may require a broader socio-technical treatment that analyzes normative expectations of integrity. We anticipate that such tensions between integrity and AI assistants may become more relevant as AI summary features become more ubiquitous and more relied upon.
@@ 9 Conclusion
AI models are being developed rapidly and integrated across devices and applications, including in contexts where they may process end-to-end encrypted or otherwise private content. In this work, we performed a detailed examination of complementary technical and legal tensions inherent to integrating AI models with E2EE applications, identifying some serious risks and concerns. Perhaps the greatest risk is in eroding E2EE as a systemic guarantee—a condition of any application or service that is held out to be E2EE—and turning it instead into merely a bonus feature that is routinely compromised in exchange for convenience.
Development of AI features in E2EE environments should be accompanied by informed multidisciplinary discussions on potential harms. Our work represents a first step towards this; we hope this conversation continues, involving free speech, privacy, and human rights advocates, civil society representatives, and others outside academia, to develop frameworks and standards for the responsible deployment of AI and E2EE. It is critical for these debates to occur promptly, before unilateral industry rollout.
@@
MK had the initial idea and co-led the overall structuring of the paper and recommendations with SP. AF, BLH, and DY contributed equally to designing and writing technical background, implications, and recommendations. DF and JL contributed equally to designing and writing legal background, implications, and recommendations. SA led the case studies and provided research support throughout other aspects of the paper. SP was faculty lead and KC provided consultation.
@@ Acknowledgments
Appendix A E2EE Features Extended Evaluation Framework
@@
Our framework distinguishes between three core axes that must be considered in the context of application features: (1) the technical implementation of the feature, and its confidentiality guarantees (or lack thereof). For considering AI assistants, we have performed this layer of analysis in Section 4.2; (2) the data inputs to the feature, i.e., what type of user data will interface with it; and (3) the level of user control over the feature. We discuss these components in more detail below.
@@
Appendix B Meta AI Disclaimers in WhatsApp
Figure 7: Meta AI’s disclaimer when typing “@MetaAI” on WhatsApp for Android the first time, as of 02/14/25.
Figure 8: Meta AI’s disclaimer when typing “@MetaAI” on WhatsApp for iOS for the first time, as of 03/15/25.
Figure 9: Meta AI’s disclaimer when direct messaging with Meta AI on WhatsApp for iOS for the first time, as of 03/15/25.